OpenStudy (anonymous):

The regular price of a child's entry ticket to a water park is $8 less than that for an adult's. The park offers half off of all entry tickets during the off-peak season. The Sandlers paid a total of $194 for 1 adult ticket and 4 child's tickets to the water park during the off-peak season. The following equation represents this situation, where x represents the regular price of an adult ticket. 194 = one-halfx + 2(x - 8) What is the regular price of a child's ticket? $68 $76 $84 $89

OpenStudy (anonymous):

194=(1/2)x+2(x-8) first, we will times the contents of the parentheses by 2 194=(1/2)x+2x-16 second, we combine like terms and add 16 to both sides 210=2(1/2)x finally, we divide both sides by 2.5 to get x by itself, and that's the answer 84=x

OpenStudy (anonymous):

wow thanks

OpenStudy (anonymous):

No problem:)

OpenStudy (anonymous):

but can you help me with 2 or 1 more please

OpenStudy (anonymous):

Sure, I can try

OpenStudy (anonymous):

x - 4y + 5z - 3x - y + 2z = ? 2x - 5y + 7z -2x - 5y - 7z -2x - 5y + 7z -2x -5y + 3z

OpenStudy (anonymous):

We need to combine all the like terms, which would be all of the xs, ys, and zs. Our final answer is -2x-5y+7z